# LMU-ASC 02/24

### v1

@@ 1 Introduction
In [11], it has been argued that the collective behavior of a tower of species is holographic and can be described in terms of species thermodynamics. This means that species carry entropy, temperature and also an overall total mass which precisely follow the rules of (black hole) thermodynamics. Indeed, the very reason why species thermodynamics can be formulated is arguably that species themselves are closely related to black holes (see also [12]). This interconnection arises due to the species scale $\Lambda_{sp}$ [13, 14, 15, 16, 17], which provides an upper bound for $\Lambda_{UV}$ in any gravitational EFT. In fact, $\Lambda_{sp}$ can be defined as (an upper bound on) the scale at which gravitational interactions become strong,
@@
up to additive log corrections. This proposal is in fact well-motivated by the observation that $L_{sp}$ corresponds to the radius of the smallest possible black hole that can be constructed as classical solution of the gravitational EFT. This is deemed a minimal black hole. Hence, the species entropy ${\cal S}_{sp}$ coincides with that of the minimal black, while the species temperature (for Schwarzschild-like black holes) is given by the inverse radius of the minimal black hole, namely
@@ 2.2 Definition of minimal black hole, species entropy and temperature
where we also provided the large-mass (equivalently low-temperature) asymptotics using that $K_{\nu}(z)\sim\sqrt{\frac{\pi}{2z}}e^{-z}$ for $z\gg 1$. The latter expression shows the expected Boltzmann suppression factor. For $q>1$ extra dimensions there is indeed a peak for an order-one value of the argument, namely at $m_{k}\sim T$. For $q=1$ extra dimensions there is no peak in the function $z^{\frac{d}{2}}K_{\frac{d}{2}}(z)$, but the expectation value of $m_{k}$ can be still estimated: for $m\ll T$, one has
@@
which for $T\lesssim\Lambda_{sp}$ is of order $(T/m)^{\gamma}$, since the integral converges even when the upper integration limit is taken to infinity. The expectation value of |$\absolutevalue{k}$ | for $q=1$ is thus a ratio of these powers, hence
@@ 2.4 Kaluza-Klein species bound states as minimal black holes
which counts how many multiparticle (species) states are inside the black hole. For a Schwarzschild-like black hole, we have then
@@
where in the last steps we wrote the asymptotic behaviour for $N\gg N_{sp}$. This regime is justified in our case due to (2.40). Compositions differing by the order of their elements are considered the same partition. Hence, to get the appropriate number of partitions we should divide by the number of compositions which are equal up to ordering. For $p=1$, this is $N_{sp}!$ and one recovers the expression used in [51]. However, for generic $p>1$ the factor $N_{sp}!$ overestimates the degeneracy. To understand this, notice that we can equivalently think of (2.36) as the number of species coming from $p$ different KK-towers. When elements of two partitions are the same but arise from different towers, we consider the partitions to be different. In other words, a multiparticle state with mass $2\Delta m_{1}$ is considered different from another with mass $\Delta m_{1}+\Delta m_{2}$, even for $\Delta m_{1}=\Delta m_{2}$. Due to this fact and assuming towers to be homogeneously populated, we estimate that the degeneracy in the number compositions with $p$ different towers is reduced from $N_{sp}!$ down to $(N_{sp}!)^{\frac{1}{p}}$. Hence, we calculate the entropy as
@@
and thus $S_{BH}\simeq N_{sp}$ as desired, even when extrapolating the result for $p\to\infty$.
@@ 3.1 Extremal and non-extremal charged black holes
Figure 1: Parameter space for Reissner–Nordström black holes. Schwarzschild black holes (blue line, ${\cal S}_{BH}=1/T_{BH}^{2}$) are at the boundary between the allowed region ($Q_{BH}^{2}>0$) and the excluded one ($Q_{BH}^{2}<0$). The red line is the line of constant extremality parameter $c=\mathcal{S}_{BH}{T}_{BH}$. For ${T}_{BH}>{T}_{0}={1/c}$, one cannot keep $c$ constant while staying in the allowed region.
@@ 3.3 Minimal charged black holes as species bound states
where one expects $a_{i}=\mathcal{O}(1)$. Thus $\alpha_{1}=\frac{M_{P}^{d}}{\Lambda^{d}}\,a_{1}$ and $\alpha_{3}=\frac{M_{P}^{2}}{\Lambda^{2}}\,a_{3}$. Therefore, away from the species limit $\alpha_{i}=\mathcal{O}(1)$ and the extremality parameter is modified by a suitable linear combination $\alpha$. In the species limit $\Lambda\ll M_{P}$ and $\alpha_{1}\gg\alpha_{3}\gg 1$ unless there is a parametric suppression of the $a_{i}$. One way to achieve this is to assume that the electromagnetic coupling $g$ is small, so that the canonical normalization in $\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ translates into factors of $g^{2}$ in the expressions for the Wilson coefficients. At any rate, insofar as the linear combination $\alpha\geq 0$ black holes can satisfy the corrected extremality bound even when formed by super-extremal constituents.
@@ 3.3.1 Non-extremal KK-like tower
Enforcing the tower weak gravity conjecture, i.e. that the entire tower be super-extremal, implies that $\beta\geq 0$. Actually, to really have a conserved $U(1)$ gauge symmetry the tower must be extremal: it should not decay and $\beta=0$. In this case, we would also have that $T_{sp}=0$. However if the $U(1)$ gauge symmetry is mildly broken, the charges $q_{n}$ are not conserved anymore and the species can decay. In this case, $\beta\neq 0$ and the species will possess a non-vanishing temperature. We will investigate the construction in string theory of this system in section 5.
@@
Then, at leading order in the species entropy the extremality parameter can be expressed as
$c\simeq\sqrt{\beta}\Lambda_{sp}^{\frac{4-d}{2}}\simeq\sqrt{\beta}({\cal S}_{sp})^{\frac{d-4}{2(d-2)}}\,,$ (3.28)
@@
$T_{sp}\simeq{\frac{c}{{\cal S}_{sp}}}={\frac{\sqrt{\beta}}{({\cal S}_{sp})^{\frac{d}{2d-4}}}}\simeq\sqrt{\beta}(\Lambda_{sp})^{d/2}\,.$ (3.29)
One can check that all thermodynamic relations for charged black holes, respectively for charged species, are satisfied. Note that, when compared to that of an uncharged tower in (2.45), the temperature derived here in the charged case is suppressed by a factor of ${\cal S}_{sp}^{-1/2}$.
@@
$\Gamma_{sp,qm}\simeq T_{sp}^{d-1}\simeq\Lambda_{sp}^{d(d-1)/2}\,.$ (3.30)
In terms of entropy suppression, this corresponds to $n=d/2$ in (3.8), i.e. $\Gamma_{sp,qm}\simeq T_{sp}{\cal S}_{sp}^{-d/2}$.
@@ 4.1 Species decay in an expanding universe
${\rm Near~{}extremal~{}Species:}\quad T_{univ}(t_{i})$ $\simeq T_{sp}^{\rm near-ext}\simeq\Lambda_{sp}^{d/2}\,.$ (4.3)
We see that initial temperature of neutral species is as high as the UV cut-off scale, whereas the initial temperature of near-extremal species is parametrically smaller by a factor $\Lambda_{sp}^{\frac{d-2}{2}}$ if $d>2$,
@@ 4.2 Species decay in the Dark Dimension scenario
${\rm Near~{}extremal~{}Species:}\quad$ $T_{sp}^{\rm near-ext}\simeq(\Lambda_{cc})^{\frac{1}{2(d-1)}}\xrightarrow[d=4]{}(\Lambda_{cc})^{\frac{1}{6}}\,.$ (4.14)
@@
This precisely agrees with the temperature (4.14) of the near-extremal species, which we argued to be the temperature and the mass scale at which species are produced during the radiation period of the expanding universe. Moreover, for phenomenological reasons it is important to point out that the charge we are considering in this scenario is associated with a new $U(1)_{\text{dd}}$ gauge symmetry possibly dynamically broken.

### v2

@@ 1 Introduction
In [11], it has been argued that the collective behavior of a tower of species is holographic and can be described in terms of species thermodynamics. This means that species carry entropy, temperature and also an overall total mass which precisely follow the rules of (black hole) thermodynamics. Indeed, the very reason why species thermodynamics can be formulated is arguably that species themselves are closely related to black holes. Indeed, in [11, 12] this connection has been outlined comparing minimal black holes to “typical” ensembles of species, whose thermodynamic quantities match. [^1] With “typical” we mean that momenta in some reference frame (corresponding to the rest frame of the associated minimal black hole) are of the order of the mass gap, while occupation numbers are order one. This interconnection arises due to the species scale $\Lambda_{sp}$ [13, 14, 15, 16, 17, 18], which provides an upper bound for $\Lambda_{UV}$ in any gravitational EFT. In fact, $\Lambda_{sp}$ can be defined as (an upper bound on) the scale at which gravitational interactions become strong,
@@
up to additive log corrections. This proposal is in fact well-motivated by the observation that $L_{sp}$ corresponds to the radius of the smallest possible black hole that can be constructed as classical solution of the gravitational EFT. This is deemed a minimal black hole. Hence, the species entropy ${\cal S}_{sp}$ coincides with that of the minimal black hole, while the species temperature (for Schwarzschild-like black holes) is given by the inverse radius of the minimal black hole, namely
@@ 2.2 Definition of minimal black hole, species entropy and temperature
where we also provided the large-mass (equivalently low-temperature) asymptotics using that $K_{\nu}(z)\sim\sqrt{\frac{\pi}{2z}}e^{-z}$ for $z\gg 1$. The latter expression shows the expected Boltzmann suppression factor. For $q>1$ extra dimensions there is indeed a peak for an order-one value of the argument, namely at $m_{k}\sim T$. For $q=1$ extra dimensions there is no peak in the function $z^{\frac{d}{2}}K_{\frac{d}{2}}(z)$, but the expectation value of $m_{k}$ can be still estimated: for $m\ll T$, the (unnormalized) expectation value of a generic power $k^{\gamma}$ is given by
@@
which for $T\lesssim\Lambda_{sp}$ is of order $(T/m)^{\gamma}$, since the integral converges even when the upper integration limit is taken to infinity. The properly normalized expectation value of |$\absolutevalue{k}$ | for $q=1$ is thus a ratio of these powers, with $\gamma=1$ and $\gamma=0$, hence
@@ 2.4 Kaluza-Klein species bound states as minimal black holes
which counts the total mass level associated to the black hole. This level can be reached combining (multi)particle states with levels up to $N_{sp}$. For a minimal Schwarzschild-like black hole, we have then
@@
where in the last steps we wrote the asymptotic behaviour for $N\gg N_{sp}$. This regime is justified in our case due to (2.40). Compositions differing by the order of their elements are considered the same partition. Hence, to get the appropriate number of partitions we should divide by the number of compositions which are equal up to ordering. For $p=1$, this is $N_{sp}!$ and one recovers the expression used in [52]. However, for generic $p>1$ the factor $N_{sp}!$ overestimates the degeneracy. To understand this, notice that we can equivalently think of (2.36) as the number of species coming from $p$ different KK-towers. When elements of two partitions are the same but arise from different towers, we consider the partitions to be different. In other words, a multiparticle state with mass $2\Delta m_{1}$ is considered different from another with mass $\Delta m_{1}+\Delta m_{2}$, even for $\Delta m_{1}=\Delta m_{2}$. Due to this fact and assuming towers to be homogeneously populated, since the quantum numbers of multiple towers multiply we estimate that the degeneracy in the number compositions with $p$ different towers is reduced from $N_{sp}!$ down to $(N_{sp}!)^{\frac{1}{p}}$. Hence, we calculate the entropy as
@@
and thus $S_{BH}\simeq N_{sp}$ as desired, even when extrapolating the result for $p\to\infty$. This combinatorial estimate agrees with the microcanonical result obtained in [12].
@@ 3.1 Extremal and non-extremal charged black holes
Figure 1: Parameter space for Reissner–Nordström black holes. Schwarzschild black holes (blue line, ${\cal S}_{BH}=1/T_{BH}^{2}$) are at the boundary between the allowed region ($Q_{BH}^{2}>0$) and the excluded one ($Q_{BH}^{2}<0$). The red line is the line of constant extremality parameter $c=\mathcal{S}_{BH}{T}_{BH}$. For ${T}_{BH}>{T}_{0}={1/c}$, one cannot keep $c$ constant while staying in the allowed region. The figure is taken from [57].
@@ 3.3 Minimal charged black holes as species bound states
where one expects $a_{i}=\mathcal{O}(1)$. Thus $\alpha_{1}=\frac{M_{P}^{d}}{\Lambda^{d}}\,a_{1}$ and $\alpha_{3}=\frac{M_{P}^{2}}{\Lambda^{2}}\,a_{3}$. Therefore, away from the species limit $\alpha_{i}=\mathcal{O}(1)$ and the extremality parameter is modified by a suitable linear combination $\alpha$. In the species limit the gravitational sector of higher-derivative corrections is expected to have $\Lambda\ll M_{P}$ and $\alpha_{1}\gg\alpha_{3}\gg 1$, unless there is a parametric suppression of the $a_{i}$, but in principle the electromagnetic sector considered above may have $\mathcal{O}(1)$ Wilson coefficients in Planck units. One way for this to occur is if the electromagnetic coupling $g$ is small, so that the canonical normalization in $\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ translates into factors of $g^{2}$ in the expressions for the Wilson coefficients [^7] A more physical intuition behind such potential discrepancy is that the species contributing to electromagnetic higher-derivative corrections can be a small, charged, subset.. At any rate, insofar as the linear combination $\alpha\geq 0$ black holes can satisfy the corrected extremality bound even when formed by super-extremal constituents, since the extremality parameter is corrected as [64]
$c^{2}=M_{BH,min}^{2}-Q_{BH,min}^{2}+2\alpha\,M_{BH,min}^{2\frac{d-4}{d-3}}\,.$ (3.24)
@@ 3.3.1 Non-extremal KK-like tower
[…]
@@
For $d>4$, in the species limit $\Lambda_{sp}\ll M_{P}$ the higher-derivative corrections to the extremality parameter scale as $\alpha\Lambda_{sp}^{2(4-d)}$ and dominate over the contributions coming from the tower, which scale as $\beta\Lambda_{sp}^{4-d}$. Then, at leading order in the species entropy the extremality parameter can be expressed as
$c\simeq\sqrt{\alpha}\Lambda_{sp}^{4-d}\simeq\sqrt{\alpha}({\cal S}_{sp})^{\frac{d-4}{d-2}}\,,$ (3.29)
@@
$T_{sp}\simeq{\frac{c}{{\cal S}_{sp}}}={\frac{\sqrt{\alpha}}{({\cal S}_{sp})^{\frac{2}{d-2}}}}\simeq\sqrt{\alpha}(\Lambda_{sp})^{2}\,.$ (3.30)
One can check that all thermodynamic relations for charged black holes, respectively for charged species, are satisfied. Note that, when compared to that of an uncharged tower in (2.45), the temperature derived here in the charged case is suppressed by a factor of ${\cal S}_{sp}^{-\frac{1}{d-2}}$.
@@
$\Gamma_{sp,qm}\simeq T_{sp}^{d-1}\simeq\Lambda_{sp}^{2(d-1)}\,.$ (3.31)
In terms of entropy suppression, this corresponds to $n=2$ in (3.8), i.e. $\Gamma_{sp,qm}\simeq T_{sp}{\cal S}_{sp}^{-2}$.
@@ 4.1 Species decay in an expanding universe
${\rm Near~{}extremal~{}Species:}\quad T_{univ}(t_{i})$ $\simeq T_{sp}^{\rm near-ext}\simeq\Lambda_{sp}^{2}\,.$ (4.3)
We see that initial temperature of neutral species is as high as the UV cut-off scale, whereas the initial temperature of near-extremal species is parametrically smaller by a factor $\Lambda_{sp}$, and
@@ 4.2 Species decay in the Dark Dimension scenario
${\rm Near~{}extremal~{}Species:}\quad$ $T_{sp}^{\rm near-ext}\simeq(\Lambda_{cc})^{\frac{2}{d(d-1)}}\xrightarrow[d=4]{}(\Lambda_{cc})^{\frac{1}{6}}\,.$ (4.14)
@@
This precisely agrees with the temperature (4.14) of the near-extremal species in $d=4$, which we argued to be the temperature and the mass scale at which species are produced during the radiation period of the expanding universe. Moreover, for phenomenological reasons it is important to point out that the charge we are considering in this scenario is associated with a new $U(1)_{\text{dd}}$ gauge symmetry possibly dynamically broken.